# Using Large Language Models for Expert Prior Elicitation in Predictive Modelling

### v1

@@ 5 Further results
Within the Appendix, we supplement our findings with further results and discussions. Firstly, we present the full prior elicitation and model selection process on a new dataset in Appendix A.5.5. Then we present additional results of prior elicitation and in-context learning (Appendix A.6, A.7, A.8), where we support the conclusions drawn in Sections 3. Next, since some of the datasets tested are in the public domain and might have appeared in the language model’s training, in Appendix A.9, we provide a discussion on whether the language model has memorised the datasets tested.
@@ 6 Conclusion
In this work, we presented a method for extracting task information from an LLM to build probabilistic linear models with fewer data points. By asking a language model to play the role of a domain expert, we were able to elicit prior distributions for predictive tasks, potentially reducing the cost of data collection and improving the accuracy of our models when we have access to few labelled examples (Section 3.1). As additionally demonstrated in Appendix A.6.2, these prior distributions can significantly impact the delivery of care in healthcare settings, where collecting labels can be costly and time-consuming, but the need for accurate predictive models is great.
We find that the prior distributions elicited from GPT-3.5-turbo achieve better accuracy or mean squared error with fewer data points than an uninformative prior for a range of tasks (Section 3.1). This performance improvement is most visible when few training examples were available, suggesting its use in environments where label collection is challenging. We also find that incorporating expert information in the task descriptions can enable further improvements in predictive accuracy, and that the language model updates its prior distributions as we would expect (Section 3.2 and Appendix A.10).
Since our approach to prior elicitation can also be seen as a direct replacement for in-context learning for datasets with numerical features, we present a framework for deciding whether, given a task, to use prior elicitation or in-context learning based on the Bayes factor (Section 3.6). However, the Bayes factor assumes that the in-context learner is performing Bayesian inference. We therefore use maximum likelihood estimation and sampling distributions to approximate the internal model being used during in-context learning and its prior and posterior distributions (Section 3.3). This allows us to inspect the truthfulness of the language model when we conduct prior elicitation (Section 3.4), as well as empirically question whether the language model is performing Bayesian inference on a given task (Section 3.5).
@@ A.1 Availability of code
Within the supplementary materials we provide code for generating all figures and results, as well as provide a pipeline for using our method on additional datasets.
@@ A.2 Extracting internal prior and posterior
Which provides us with the sampling distribution:
@@
With this, we can generate a sampling distribution using $\{\varphi_{\text{MLE}}^{(s)}\}_{s=0}^{s=S}$ that approximates $p_{M,T}(\varphi)$ as $S\rightarrow\infty$. Additionally, by further incorporating the training dataset $\mathcal{D}$:
@@ A.3 Does the language model use the model class we specify?
As part of comparing prior elicitation to in-context learning, we use maximum likelihood sampling to extract the internal parameter distributions (discussed in Section 3.3), which we use to test whether the language model is: (1) providing a similar prior distribution through elicitation to the one it is using in-context, and (2) empirically test whether the language model is approximating Bayesian inference. However, when extracting the internal parameters of the in-context model, we assume that the language model is applying the model class that we specify in the task description.
@@
In Figure 7 we show this true $\hat{y}_{n}$ against the in-context prediction $y_{n}$. Interestingly, even given the simple linear equation we ask the language model to mimic, we find that it struggles to provide good predictions in-context. This suggests that the language model is not always using the model class for in-context predictions that we specify in the task description, indicating that in its current state, in-context learning might be inadequate for predictive tasks with primarily numerical features. This figure also shows that the language model was hesitant to predict negative values and preferred positive predictions – a strange behaviour that clearly restricts its use as an emulator of a linear model.
However, for all other prior extraction tasks presented in Section 3.3, we find that the language model uses a linear model in-context, demonstrated by the low values of MSE loss between the predictions of our MLE samples and the true in-context predictions, shown in Figure 11. This is often not the case when the language model is presented with demonstration samples, as in the posterior extraction presented in Figure 5, in which the language model uses a linear combination of features to predict the targets for Heart Disease and Wine Quality only.
@@ A.4 Energy distance and statistic
The energy statistic [33], which is based on the energy distance metric [50], is a method for measuring the similarity of two probability distributions. Importantly, it allows us to test whether two $N$-dimensional probability distributions are equal.
@@
Where ${M_{XY}}_{i,j}=||x_{i}-y_{j}||_{2}$ is the pairwise distance matrix.
From this, we construct the energy statistic as:
@@
This statistic takes values between 00 and $1$, and tests whether two statistical distributions are equal.
@@ A.5.1 Dataset details
Within this work we evaluate our method on $6$ datasets; $1$ of which is synthetic, $1$ is privately collected, and $4$ are publicly accessible. This allows us to provide both reproducible experiments and test the language model’s ability to suggest a good prior on datasets not seen during training.
@@ A.5.2 Evaluating prior elicitation
For the UTI dataset, we split the dataset such that each participant’s data is only in either the testing or training set, as some participants produced multiple examples. For the other classification tasks, splits are stratified using the targets. To test the posterior performance after observing the given number of data points in the experiments, we randomly select examples from the training set.
@@
We then construct a logistic or linear regression model, where we set the prior distribution over the bias term as a mixture of $100$ distributions of $\theta_{0}\sim\mathcal{N}(0,1)$. In the case of linear regression, we also use a noise term $\epsilon\sim\text{Half-Cauchy}(\beta=1)$, with a Half-Cauchy prior.
In the uninformative case, we set $\theta\sim\mathcal{N}(0,1)$ for each feature.
@@
To produce the results in Figure 2(a) and 8, we then sample the posterior predictive distribution over our test set and calculate the accuracy or mean squared error for each posterior parameter sample. These metric values are then presented in Figure 2(a) and 8, which show their histograms. To produce Figure 2(b) and 13, we also calculate the mean accuracy and mean squared error for each of the test splits before plotting their mean and standard deviation.
@@ A.5.3 Extracting the in-context model’s hidden prior and posterior
To further study in-context learning, we used maximum likelihood sampling to extract approximate samples of the internal model’s prior and posterior using the method described in Section 2.3.2. For each of the tasks studied, we randomly generate $25$ feature values from $X\sim\mathcal{U}[-5,5]$ $5$ times over and ask the language model to predict their regression label or positive classification probability based on the task description given and with the model class specified (in our case a linear or logistic regressor). This is done for each of the $100$ task descriptions we have for the dataset. We then fit a linear regressor to the predicted regression labels or the classification logits  [^3] Calculated using the inverse logistic function: $\text{logits}=\ln p/(1-p)$ of the predicted probabilities, using maximum likelihood estimation. The parameters of this linear model are then taken as a single sample of the approximated in-context predictive model.
@@
We can then measure, as in Section 3.4, whether the parameter samples are similar to those sampled from the elicited distributions. This is done by sampling $10000$ parameter values from the elicited distributions and calculating the energy statistic (Appendix A.4) between this sample and the extracted in-context sample. This is done to test whether the language model is using the same distribution we elicit from it, as the one it uses in-context. The results of this are presented in Table 1.
@@
The results presented in Table 2 show the result of this experiment.
@@ A.5.4 Calculating Bayes factor
As part of our solution for deciding whether to use prior elicitation or in-context learning for a given dataset, we calculate the Bayes factor (presented in Section 2.4).
@@
For both sets of prior predictive samples, we calculate their log-likelihood using either cross-entropy loss or mean squared error, depending on whether the task is classification or regression. For each dataset, for each test split, for each method, we calculate the mean of this log-likelihood over the samples. With these values, for each dataset and test split, we measure the difference between the value for the elicited prior predictions and the in-context predictions to get the log Bayes factor.
We then report the mean Bayes factor over the splits along with the standard deviation in Figure 6, which helps us to decide, given a dataset, which method is most appropriate.
@@ A.5.5 Given a new dataset
    With the randomly generated feature values and the in-context predictions, use maximum likelihood estimation to estimate the underlying linear model for each task description. For the classification case, this is more easily done if the language model is asked to provide the probability of a positive prediction rather than the prediction itself.
@@
    Use a method to calculate the the difference between two samples of a distribution to understand the difference between the posterior samples calculated through Monte Carlo sampling of the in-context prior and the samples of the in-context posterior.
@@ A.6 Further prior elicitation results
Here, we present the results of the same experiment as discussed in Section 3.1, applied to two further datasets discussed in Section 2.5.
The results of a logistic regression model on a test set after observing training data and using a prior collected through elicitation of the language model or an uninformative prior are shown in Figure 8.
@@
Further, for the Heart Disease dataset, we find that the prior elicited from the language model performed comparably to the uninformative prior. This is because the language model provided mean = 0 and std = 1 for a large number of priors elicited, suggesting that the language model was unsure about the prior distribution on the parameters for a model predicting heart disease. The elicited parameter distributions can be seen in Figure 11, which contextualises these results.
@@ A.6.2 UTI accuracy with the label collection context
Since we have access to the underlying study that collected the UTI dataset, we are able to understand in more detail how the accuracy of the predictive model varies in time, as well as with the number of data points collected.
In Figure 10 we plot the length of time it took to collect the given number of labels tested in the experiments in Section 3.1 against the accuracy of a model using an uninformative prior and one elicited from a language model. This result illustrates the usefulness of prior elicitation, since we are able to achieve comparable (and ultimately better) accuracy with less data points, and therefore less data collection time.
@@
It is important to note, that the accuracy in Figure 10 is not achieved as a result of training on the earliest collected data points. This is because in our analysis, we perform cross-validation and use a subset of the data from later time period (Section A.5.1). There were changes to the in-home devices during the study, such as the addition of a device for collecting sleep information, which we use within the dataset evaluated in the main results. However, if all of the data collection devices had been used from the beginning of the study, we might expect the accuracy in time to resemble the figure presented here.
@@ A.7 Is the language model being truthful when providing priors through elicitation?
Figure 11 visualises the different prior parameter distributions for the UTI, Breast Cancer, California Housing, Wine Quality, and Heart disease datasets in each feature dimension. Of note is the large differences between the elicited prior distributions and the ones used when performing in-context learning, which supports the findings in Section 3.4. In the in-context case, the prior distributions frequently have their modes close to 00, suggesting the internal model is unsure about the relationship between features and targets. On the other hand, the elicited priors often represent a stronger belief about the correlation between features and targets, with most of the histograms achieving modes greater than or less than 00.
@@
Further, the prior distribution on the parameters for the Heart Disease dataset have more variation in their values in comparison to that extracted from the language model’s internal model. This tells us that the elicited prior parameters represent more uncertainty in their value and are therefore more flexible when calculating the posterior on observed data.
@@ A.8 Testing empirically whether the language model is approximating Bayesian inference
On the California Housing dataset, the posteriors do not match as closely. The histograms given in Figure 12, along with the comparatively larger energy distance, show a significant difference in the distributions. In the extracted posterior, almost all of the values are contained within a small range around zero, whilst the posterior sampled from MCMC methods shows a larger variation in possible values. Additionally, in this cases, the energy statistic is far greater than for other datasets and suggests that for the California Housing regression task, in-context learning is not approximating Bayesian inference.
@@ A.9 Dataset memorisation
To better contextualise the use of in-context learning on the datasets tested, we will utilise the method proposed in Bordt et al. [42] to quantify the extent to which the language model already knows information about the dataset.
Table 3 shows the results of the header completion test described in Bordt et al. [42], which tests whether the language model has memorised the first few rows of the dataset. The column names, along with a given number of the first rows are given to the language model as a string, with the final of these rows truncated. The language model is then asked to generate the next $500$ tokens.
We use the Levenshtein distance [51, 52], a string matching metric, to calculate the agreement between the true header completion and the one provided by the language model. The Levenshtein distance counts the number of single-character changes required to transform one string into another, only allowing for insertions, deletions, and substitutions. The value can take a minimum value of 00 – suggesting the strings are equal and a maximum value of the greatest length of the two strings.
@@
The values in Table 3 demonstrate that the language model has memorised the first few rows of the Breast Cancer and Heart Disease datasets completely, demonstrated by a normalised Levenshtein distance value of 0. This could be because the first few rows of these datasets commonly appear within the language model’s training data. For California Housing and Wine Quality, we see some level of memorisation, with the language model being able to reproduce a large amount of the true header, however, to a much lesser extent than what we see for Breast Cancer and Heart Disease. As we might expect, when we perform the header test on the synthetic dataset generated using the relationship $y=2x_{1}-x_{2}+x_{3}$, we get a Normalised Levenshtein distance of $0.7$, which shows a large difference between the true header string and the language models completion. The UTI dataset has not been tested, as it contains sensitive information and is currently a private dataset.
@@
We supplement these results with those in Table 4, which show the results of the row completion test. Here we calculate the normalised Levenshtein distance of the true completions and the language model’s completion of a random row from the dataset. In these tests, the language model is provided with $10$ rows (in order from the original dataset) and is asked to complete the following row. The values in Table 4 represent the mean and standard deviation of the test using $25$ random rows. Here, in contrast to the header test, we see that the language model has not memorised all rows from the dataset. This is because the language model’s completion differs from the true row completion in all cases to varying degrees.

### v2

@@ 5 Further results
In the Appendix, we supplement our findings with further results and discussions. Firstly, we present the complete prior elicitation and model selection process on a new dataset in Appendix A.5.5. Then we present additional results of prior elicitation and in-context learning (Appendix A.6, A.7, A.8), where we support the conclusions drawn in Sections 3. Next, since some of the datasets tested are in the public domain and might have appeared in the language model’s training, in Appendix A.9, we provide a discussion on whether the language model has memorised the datasets tested.
@@ 6 Conclusion
In this work, we presented a method for extracting task information from an LLM to build probabilistic linear models with fewer data points. By asking a language model to play the role of a domain expert, we were able to elicit prior distributions for predictive tasks, potentially reducing the cost of data collection and improving the accuracy of our models when we have access to few labelled examples (Section 3.1). As also demonstrated in Appendix A.6.2, these prior distributions can significantly impact the delivery of care in healthcare settings, where collecting labels can be costly and time-consuming, but the need for accurate predictive models is great.
We find that the prior distributions elicited from GPT-3.5-turbo achieve better accuracy or mean squared error with fewer data points than an uninformative prior for a variety of tasks (Section 3.1). This performance improvement is most visible when few training examples were available, suggesting its use in environments where label collection is challenging. We also find that incorporating expert information into the task descriptions can enable further improvements in predictive accuracy, and that the language model updates its prior distributions as we would expect (Section 3.2 and Appendix A.10).
Since our approach to prior elicitation can also be seen as a direct replacement for in-context learning for datasets with numerical features, we present a framework for deciding whether, given a task, to use prior elicitation or in-context learning based on the Bayes factor (Section 3.6). However, the Bayes factor assumes that the in-context learner is performing Bayesian inference. Therefore, we use maximum likelihood estimation and sampling distributions to approximate the internal model used during in-context learning and its prior and posterior distributions (Section 3.3). This allows us to inspect the truthfulness of the language model when we conduct prior elicitation (Section 3.4), as well as empirically question whether the language model is performing Bayesian inference on a given task (Section 3.5).
@@ A.1 Availability of code
Within the supplementary materials, we provide code for generating all figures and results, as well as provide a pipeline for using our method on additional datasets.
@@ A.2 Extracting internal prior and posterior
This provides us with the sampling distribution:
@@
With this, we can generate a sampling distribution using $\{\varphi_{\text{MLE}}^{(s)}\}_{s=0}^{s=S}$ that approximates $p_{M,T}(\varphi)$ as $S\rightarrow\infty$. Furthermore, by further incorporating the training dataset $\mathcal{D}$:
@@ A.3 Does the language model use the model class we specify?
As part of comparing prior elicitation with in-context learning, we use maximum likelihood sampling to extract the internal parameter distributions (discussed in Section 3.3), which we use to test whether the language model is: (1) providing a similar prior distribution through elicitation to the one it is using in-context, and (2) empirically test whether the language model is approximating Bayesian inference. However, when extracting the internal parameters of the in-context model, we assume that the language model is applying the model class that we specify in the task description.
@@
In Figure 7 we show the true $\hat{y}_{n}$ against the in-context predictions $y_{n}$. Interestingly, even given the simple linear equation we ask the language model to mimic, we find that it struggles to provide good predictions in-context. This suggests that the language model is not always using the model class for in-context predictions that we specify in the task description, indicating that in its current state, in-context learning might be inadequate for predictive tasks with primarily numerical features. This figure also shows that the language model was hesitant to predict negative values and preferred positive predictions – a strange behaviour that clearly restricts its use as an emulator of a linear model.
However, for all other prior extraction tasks presented in Section 3.3, we find that the language model uses a linear model in-context, demonstrated by the low MSE loss values between the predictions of our MLE samples and the true in-context predictions, shown in Figure 11. This is often not the case when the language model is presented with demonstration samples, as in the posterior extraction presented in Figure 5, in which the language model uses a linear combination of features to predict the targets for Heart Disease and Wine Quality only.
@@ A.4 Energy distance and statistic
The energy statistic [33], which is based on the energy distance metric [50], is a method to measure the similarity of two probability distributions. Importantly, it allows us to test whether two $N$-dimensional probability distributions are equal.
@@
Here ${M_{XY}}_{i,j}=||x_{i}-y_{j}||_{2}$ is the pairwise distance matrix.
From this, we construct the energy statistic as follows:
@@
This statistic takes values between 00 and $1$ and tests whether two statistical distributions are equal.
@@ A.5.1 Dataset details
In this work, we evaluate our method on $6$ datasets; $1$ of which is synthetic, $1$ is privately collected, and $4$ are publicly accessible. This allows us to provide both reproducible experiments and test the language model’s ability to suggest a good prior on datasets not seen during training.
@@ A.5.2 Evaluating prior elicitation
For the UTI task, we split the dataset so that each participant’s data is only in either the testing or training set, as some participants produced multiple examples. For the other classification tasks, the splits are stratified using the targets. To test the posterior performance after observing the given number of data points in the experiments, we randomly select examples from the training set.
@@
We then construct a logistic or linear regression model, where we set the prior distribution over the bias term as a mixture of $100$ distributions of $\theta_{0}\sim\mathcal{N}(0,1)$. In the case of linear regression, we also use a noise term $\epsilon\sim\text{Half-Cauchy}(\beta=1)$, with a Half-Cauchy prior. In the uninformative case, we set $\theta\sim\mathcal{N}(0,1)$ for each feature.
@@
To produce the results in Figures 2(a) and 8, we then sample the posterior predictive distribution over our test set and calculate the accuracy or mean squared error for each posterior parameter sample. These metric values are then presented in Figures 2(a) and 8, which show their histograms. To produce Figures 2(b) and 13, we also calculate the mean accuracy and the mean squared error for each of the test splits before plotting their mean and standard deviation.
@@ A.5.3 Extracting the in-context model’s hidden prior and posterior
To further study in-context learning, we used maximum likelihood sampling to extract approximate samples of the internal model’s prior and posterior using the method described in Section 2.3.2. For each of the tasks studied, we randomly generate $25$ feature values from $X\sim\mathcal{U}[-5,5]$ $5$ times and ask the language model to predict their regression label or positive classification probability based on the task description given and with the specified model class (in our case a linear or logistic regressor). This is done for each of the $100$ task descriptions that we have for the dataset. We then fit a linear regressor to the predicted regression labels or the classification logits  [^3] Calculated using the inverse logistic function: $\text{logits}=\ln p/(1-p)$ of the predicted probabilities, using maximum likelihood estimation. The parameters of this linear model are then taken as a single sample of the approximated in-context predictive model.
@@
We can then measure, as in Section 3.4, whether the parameter samples are similar to those sampled from the elicited distributions. This is done by sampling $10000$ parameter values from the elicited distributions and calculating the energy statistic (Appendix A.4) between this sample and the extracted in-context sample. This is done to test whether the language model is using the same distribution we elicit from it as the one it uses in-context. The results of this are presented in Table 1.
@@
The values presented in Table 2 show the result of this experiment.
@@ A.5.4 Calculating Bayes factor
As part of our solution to decide whether to use prior elicitation or in-context learning for a given dataset, we calculate the Bayes factor (presented in Section 2.4).
@@
For both sets of prior predictive samples, we calculate their log-likelihood using either cross-entropy loss or mean squared error, depending on whether the task is classification or regression. For each dataset, for each test split, and for each method, we calculate the mean of this log-likelihood over the samples. With these values, for each dataset and test split, we measure the difference between the value for the elicited prior predictions and the in-context predictions to get the log Bayes factor.
We then report the mean Bayes factor over the splits along with the standard deviation in Figure 6, which helps us to decide, given a dataset, which method is the most appropriate.
@@ A.5.5 Given a new dataset
    With randomly generated feature values and in-context predictions, use maximum likelihood estimation to estimate the underlying linear model for each task description. For the classification case, this is done more easily if the language model is asked to provide the probability of a positive prediction rather than the prediction itself.
@@
    Use a method to calculate the difference between two samples of a distribution to understand the difference between the posterior samples calculated through Monte Carlo sampling of the in-context prior and the samples of the in-context posterior.
@@ A.6 Further prior elicitation results
Here, we present the results of the same experiment as discussed in Section 3.1, applied to two other datasets discussed in Section 2.5.
The results of a logistic regression model on a test set after observing training data and using a prior elicited from the language model or using an uninformative prior are shown in Figure 8.
@@
Furthermore, for the Heart Disease dataset, we find that the prior elicited from the language model performed comparably to the uninformative prior. This is because the language model provided mean = 0 and std = 1 for a large number of priors elicited, suggesting that the language model was unsure about the prior distribution on the parameters for a model predicting heart disease. The parameter distributions elicited can be seen in Figure 11, which contextualises these results.
@@ A.6.2 UTI accuracy with the label collection context
Since we have access to the underlying study that collected the UTI dataset, we can understand in more detail how the accuracy of the predictive model varies in time, as well as with the number of data points collected.
In Figure 10 we plot the time it took to collect the given number of labels tested in the experiments in Section 3.1 against the accuracy of a model using an uninformative prior and one elicited from a language model. This result illustrates the usefulness of prior elicitation, since we are able to achieve comparable (and ultimately better) accuracy with fewer data points, and therefore less data collection time.
@@
It is important to note that the accuracy in Figure 10 is not achieved as a result of training on the earliest collected data points. This is because in our analysis we perform cross-validation and use a subset of the data from a later time period (Section A.5.1). There were changes to the in-home devices during the study, such as the addition of a device to collect sleep information, which we use within the dataset evaluated in the main results. However, if all data collection devices had been used from the beginning of the study, we could expect the accuracy in time to resemble the figure presented here.
@@ A.7 Is the language model being truthful when providing priors through elicitation?
Figure 11 visualises the different prior parameter distributions for the UTI, Breast Cancer, California Housing, Wine Quality, and Heart Disease datasets in each feature dimension. Notable are the large differences between the elicited prior distributions and the ones used when performing in-context learning, which supports the findings in Section 3.4. In the in-context case, prior distributions frequently have modes close to 00, suggesting that the internal model is not sure about the relationship between features and targets. However, the elicited priors often represent a stronger belief in the correlation between features and targets, with most histograms achieving modes greater than or less than 00.
@@
Furthermore, the prior distribution on the parameters for the Heart Disease dataset have more variation in their values in comparison to that extracted from the language model’s internal model. This tells us that the elicited prior parameters represent more uncertainty in their value and are therefore more flexible when calculating the posterior on observed data.
@@ A.8 Testing empirically whether the language model is approximating Bayesian inference
On the California Housing dataset, the posteriors do not match as closely. The histograms given in Figure 12, together with the comparatively larger energy distance, show a significant difference in distributions. In the extracted posterior, almost all of the values are contained within a small range around zero, whilst the posterior sampled from MCMC methods shows a larger variation in possible values. Additionally, in these cases, the energy statistic is far greater than for other datasets and suggests that, for the California Housing regression task, in-context learning is not approximating Bayesian inference.
@@ A.9 Dataset memorisation
To better contextualise the use of in-context learning in the tested datasets, we will use the method proposed in Bordt et al. [42] to quantify the extent to which the language model already knows information about the dataset.
Table 3 shows the results of the header completion test described in Bordt et al. [42], which tests whether the language model has memorised the first few rows of the dataset. The column names along with a given number of the first rows are given to the language model as a string, with the final of these rows truncated. The language model is then asked to generate the next $500$ tokens.
We use the Levenshtein distance [51, 52], a string matching metric, to calculate the agreement between the true header completion and the one provided by the language model. The Levenshtein distance counts the number of single-character changes required to transform one string into another, only allowing for insertions, deletions, and substitutions. The value can take a minimum value of 00 – suggesting that the strings are equal and a maximum value of the greatest length of the two strings.
@@
The values in Table 3 demonstrate that the language model has completely memorised the first few rows of the Breast Cancer and Heart Disease datasets, demonstrated by a normalised Levenshtein distance value of 0. This could be because the first few rows of these datasets commonly appear within the language model’s training data. For California Housing and Wine Quality, we see some level of memorisation, with the language model being able to reproduce a large amount of the true header, however, to a much lesser extent than what we see for Breast Cancer and Heart Disease. As we might expect, when we perform the header test on the synthetic dataset generated using the relationship $y=2x_{1}-x_{2}+x_{3}$, we get a normalised Levenshtein distance of $0.7$, which shows a large difference between the true header string and the completion from the language model. The UTI task has not been tested, as it contains sensitive information and is currently a private dataset.
@@
We supplement these results with those of Table 4, which shows the results of the row completion test. Here we calculate the normalised Levenshtein distance of the true completions and the language model’s completion of a random row from the dataset. In these tests, the language model is provided with $10$ rows (in order from the original dataset) and is asked to complete the following row. The values in Table 4 represent the mean and standard deviation of the test using $25$ random rows. Here, in contrast to the header test, we see that the language model has not memorised all rows from the dataset. This is because the language model’s completion differs from the true row completion in all cases to varying degrees.